OpenStudy (anonymous):

A horse is extremely thirsty and is drinking out of a cylindrical trough that has a 3 meter diameter and a height of 1 meter, so if the horse is drinking 1/2 m^3/min, how fast is the water level decreasing when it is down to 1/3m?

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

wow this is a bit hard what grade is this becuase im just in 8th

OpenStudy (anonymous):

area is (3/2)^2 ft^2

OpenStudy (anonymous):

in fraction?

OpenStudy (anonymous):

oh i mean meters

OpenStudy (anonymous):

ah, I see.. so how would i get it in meters?

OpenStudy (anonymous):

fractions*

OpenStudy (bibby):

use unit cancellation. multiply by the foot/meter conversion twice

OpenStudy (anonymous):

\[\large (\frac{ 3 }{ 2 })^2=\frac{ 9 }{ 4}m^2\]

OpenStudy (bibby):

Actually there are no feet here. it's all in meters :V

OpenStudy (anonymous):

so, the fraction has gotta be in pi, no?

OpenStudy (anonymous):

\[\huge \frac{\frac{ 1 }{ 2 }m^3/\min}{\frac{ 9 }{ 4 }m^2}=\]

OpenStudy (anonymous):

oh yes i forgot pi

OpenStudy (anonymous):

good call

OpenStudy (anonymous):

1/9pi right?

OpenStudy (anonymous):

\[\huge \frac{\frac{ 1 }{ 2 }m^3/\min}{\frac{ 9\pi^2 }{ 4 }m^2}=\]

OpenStudy (anonymous):

2/81pi^2 right?

OpenStudy (anonymous):

\[\huge \frac{\frac{ 1 }{ 2 }m^3}{\frac{ 9\pi^2 }{ 4 }m^2\min}=\frac{ \frac{ 1 }{ 2 } }{ \frac{ 9\pi^2 }{ 4 } }m/\min\]

OpenStudy (anonymous):

ya, the answer above is what I'm getting, but it cant be right because of the square

OpenStudy (anonymous):

=\[\Large \frac{ 2 }{ 9\pi^2 } m/\min\]

OpenStudy (anonymous):

so, that the final or do i keep solving?

OpenStudy (anonymous):

that should be it

OpenStudy (anonymous):

it's not that because there can't be a squared

OpenStudy (anonymous):

that's what it says on my paper anyways

OpenStudy (anonymous):

oh you're right the pi isn't squared

OpenStudy (anonymous):

dunno how i messed that up, but the rest is still correct

OpenStudy (anonymous):

alrght, so how can we determine final?

OpenStudy (anonymous):

i might have screwed up too by accident

OpenStudy (anonymous):

\[\Large \frac{ 2 }{ 9\pi } m/\min\]

OpenStudy (anonymous):

thanks again pal